# World-Env: Leveraging World Model as a Virtual Environment for VLA Post-Training

### v3

@@ B.1 Deatils of Scale Head
$\mathcal{L}_{\text{NLL}}=\underbrace{|\mathbf{a}_{\text{gt}}-\bm{\mu}|\cdot e^{-\bm{\beta}}}_{\text{Data fit}}+\underbrace{\bm{\beta}}_{\text{Uncertainty penalty}}+\log 2.$ (8)
@@
B.2 Deatils of World Simulator
We adopt the original implementation of the EVAC world model (Jiang et al., 2025b) and retain its training configuration. We show an overview in Figure 8. The generation process starts from a reference image, whose CLIP (Radford et al., 2021) features provide style guidance. This signal is integrated into the diffusion model via cross-attention. Action information is encoded as a spatial action map and concatenated with visual features at the feature level. The fused representation drives the diffusion network to generate future frames through iterative denoising, followed by a video decoder to produce the final output. The EVAC model was originally designed for dual-arm robotic platforms with 14-dimensional (14D) action vectors (7D per arm). In contrast, the LIBERO benchmark employs a single-arm robot with 7D actions (6D end-effector pose + 1D gripper state). To maintain compatibility with the EVAC architecture, we zero-pad the unused 7D action dimensions during training, preserving the input interface while adapting to the target hardware.
B.3 Details of Reward Head
Our VLM-guided instant reflector integrates a pretrained vision-language model (LLaVA (Liu et al., 2023c)) with a lightweight reward head that predicts continuous reward signals, see Figure 9 for an overview. The VLM backbone is kept frozen to preserve its semantic capabilities, and only the reward head is trained. Given a video sequence $\{f_{1},\dots,f_{N}\}$ generated by the world simulator, we uniformly sample 32 frames as visual input. The language prompt is formatted as: “Watch the video and determine whether it completes the task: $\{\mathbf{g}\}$ — answer only ‘Yes’ or ‘No’.” The VLM processes this input and extracts a pooled embedding, which is projected by the reward head to a scalar. A sigmoid activation yields a continuous reward $R\in[0,1]$, interpreted as the task completion probability. The reward head is trained with binary cross-entropy loss, using a batch size of 8, learning rate $1\times 10^{-4}$, Adam optimizer, and 50 epochs, with input frames center-cropped to $\times384$ resolution.
Appendix C Analysis of World Simulator
@@
Figure 8: Overview of the world simulator.
Figure 9: Network architecture of instance reflector.
C.1 Data Analysis and Distribution
@@
C.2 More Results of World Simulator
Figures 11 and 12 show additional trajectories generated by the world simulator, demonstrating its ability to synthesize both successful and failed task executions.

### v5

@@ 9.1 Deatils of Scale Head
$\mathcal{L}_{\text{NLL}}=\underbrace{|\mathbf{a}_{\text{gt}}-\boldsymbol{\mu}|\cdot e^{-\boldsymbol{\beta}}}_{\text{Data fit}}+\underbrace{\boldsymbol{\beta}}_{\text{Uncertainty penalty}}+\log 2.$ (9)
@@
9.2 Details of Reward Head
Our VLM-guided instant reflector integrates a pretrained vision-language model (LLaVA [35]) with a lightweight reward head that predicts continuous reward signals, see Figure 10 for an overview. The VLM backbone is kept frozen to preserve its semantic capabilities, and only the reward head is trained. Given a video sequence $\{f_{1},\dots,f_{N}\}$ generated by the world simulator, we uniformly sample 32 frames as visual input. The language prompt is formatted as: “Watch the video and determine whether it completes the task: $\{\mathbf{g}\}$ — answer only ‘Yes’ or ‘No’.” The VLM processes this input and extracts a pooled embedding, which is projected by the reward head to a scalar. A sigmoid activation yields a continuous reward $R\in[0,1]$, interpreted as the task completion probability. The reward head is trained with binary cross-entropy loss, using a batch size of 8, learning rate $1\times 10^{-4}$, Adam optimizer, and 50 epochs, with input frames center-cropped to $\times384$ resolution.
@@ 9.2 Details of Reward Head
Figure 10: Network architecture of instant reflector.
Table 7: Performance of the instant reflector under different $\eta$. The selected value $\eta=0.5$ achieves a balanced trade-off across accuracy, precision, recall, and F1-score.
$\eta$ Accuracy Precision Recall F1-score
0.2 0.825 0.667 1.0 0.8
0.4 0.875 0.764 0.928 0.838
0.5 (Ours) 0.925 0.867 0.928 0.896
0.6 0.875 0.846 0.785 0.815
0.8 0.85 0.9 0.642 0.75
10 Analysis of Instant Reflector
To investigate the sensitivity of model performance to the hyperparameter $\eta$, we conduct a series of experiments by varying $\eta$ over a predefined range (e.g., 0.2, 0.4, 0.5, 0.6, 0.8). As shown in Table 7, we report four standard evaluation metrics: accuracy, precision, recall and F1-score. The results indicate that $\eta=0.5$ yields consistently strong performance across all metrics.
11 Analysis of World Simulator
11.1 Data Analysis and Distribution
@@
11.2 Ablation Studies
We evaluate the impact of training data, our proposed geometry-aware VGGT feature injection, and alternative representations (such as SAM and DINO) on the performance of the world simulator. Quantitative results are presented in Table 9, where “w/o extra” denotes the model trained without additional data, and “w/o VGGT” refers to the variant without our geometry-aware feature injection strategy. We evaluate using standard metrics: FID [19], FVD [48], PSNR [23], SSIM [52], and LPIPS [59]. As shown in Table 9, both the expanded training data and the VGGT latent features significantly contribute to building a more robust and physically consistent world model. Additional, we perturb the world model’s outputs with: (i) Gaussian noise (mean = 0, variance = 0.1), and (ii) color perturbation, randomly adjusting brightness ($\pm 20\%$), contrast ($\pm 20\%$), saturation ($\pm 20\%$), and hue ($\pm 0.1$). As shown in Table 8, thanks to the powerful pretrained VLM and our data augmentation, these perturbations only incur minor performance drop, validating our robustness to imperfect early-stage predictions.
Task Gaussian noise Color perturbation Original (Ours)
LIBERO-Spatial 85.4 87.0 87.6
Table 8: Impact of world model perturbations on final results
Table 9: Quantitative comparison of world models.
[…]
w/o extra 100.9978 116.7792 20.1600 0.7450 0.1385
w/o VGGT 70.9232 114.5264 23.1841 0.8219 0.0806
Ours with DINO 41.7398 74.9209 23.4069 0.8511 0.0602
Ours with SAM 40.6276 74.0298 23.6022 0.8532 0.0599
Ours (CLIP+VGGT) 39.1941 73.5313 23.8343 0.8579 0.0562
@@ 11.2 Ablation Studies
11.3 More Results
Figures 12 and 13 show additional trajectories generated by the world simulator, demonstrating its ability to synthesize both successful and failed task executions.